OpenStudy (anonymous):

(the question is big but the answer is easy! PLEASE HELP!!!!) the type of a software program was $50.00.at one store,the program was on sale for 10% off the retail price.jason had a coupon for an additional 10% off the sale price.at another store,gabrielle bught the same program for 20% off the retail price.did the 2 people pay the same price?

OpenStudy (anonymous):

do anyone know the answer?????

OpenStudy (anonymous):

SAT huh. Yes

OpenStudy (anonymous):

Yes they payed the same price.

OpenStudy (anonymous):

is it like 50*20%=$10 50-10=$40 and 50*20%=10 50-10=$40

OpenStudy (anonymous):

yes. It is.

OpenStudy (anonymous):

nope..... just middle school question!!!!!!!!!!!!!